OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-3, -15) and is perpendicular to -6x + 8y = 3 Please show all of your work.

OpenStudy (amistre64):

-6x + 8y = 3 ; a perp line is simply swap cooefs and switch operators, and trash the constant 8x + 6y = C ; recalibrate with the given point for C -3 -15 -24 -90 = -114 8x +6y = -114 ; then rearrange into slope,int format

OpenStudy (anonymous):

thanks